### v1

# POV Learning: Individual Alignment of Multimodal Models using Human Perception

Abstract
Aligning machine learning systems with human expectations is mostly attempted by training with manually vetted human behavioral samples, typically explicit feedback. This is done on a population level since the context that is capturing the subjective Point-Of-View (POV) of a concrete person in a specific situational context is not retained in the data.
However, we argue that alignment on an individual level can boost the subjective predictive performance for the individual user interacting with the system considerably. Since perception differs for each person, the same situation is observed differently. Consequently, the basis for decision making and the subsequent reasoning processes and observable reactions differ.
We hypothesize that individual perception patterns can be used for improving the alignment on an individual level. We test this, by integrating perception information into machine learning systems and measuring their predictive performance wrt. individual subjective assessments. For our empirical study, we collect a novel data set of multimodal stimuli and corresponding eye tracking sequences for the novel task of Perception-Guided Crossmodal Entailment and tackle it with our Perception-Guided Multimodal Transformer.
Our findings suggest that exploiting individual perception signals for the machine learning of subjective human assessments provides a valuable cue for individual alignment. It does not only improve the overall predictive performance from the point-of-view of the individual user but might also contribute to steering AI systems towards every person’s individual expectations and values.
1 Motivation
Since the release of GPT4 [15] multimodal transformers have received a great amount of public attention.
They show impressive capabilities on tasks like visual question answering that take image and text as input and generated text as output. However, aligning machine learning models with an individual person’s point-of-view (POV) and expectations has not attracted much research yet.
To improve the alignment of machine learning models according to individual expectations and preferences we first have to capture personality cues that are potential indicators for personal preferences and values. For a stimulus that consist of text and images the visual perception process is a potential signal that indicates an individual’s evaluation of said stimulus (see Fig. 1).
Figure 1: Visualisation of human visual perception when given the task: Are the central objects in the image mentioned in the caption? The red areas are showing the focus of human attention.
The other required component is a machine learning objective that is an indicator for individual alignment. When accounting for tasks that measure individual behavior, a key feature is that the same stimuli might be differently assessed by different individuals. Contrary to traditional machine learning tasks, this is not regarded as noise or an undesirable low inter-rater reliability in POV learning. Instead we are aiming for individual alignment, so individually differing assessments of the same stimuli is a key feature to be exploited.
In summary, a POV machine learning task as illustrated in Fig. 2 is about maximizing the POV-specific objective (here, crossmodel entailment), given the stimulus (here, text+image pairs) plus the individual alignment signal (here, eye-tracking perception traces).
[…]
Figure 2: Illustration of a POV machine learning task: Given the stimulus and the individual alignment signals, a machine learning model is conditioned on the POV-specific objective to predict the subjective inference of an individual.
Based on that setup, we investigate how alignment of multimodal models on an individual level can be improved by exploiting human visual perception traces. Our contributions can be summarized as follows:
  1. 1.
    For measuring individual alignment, we propose the novel POV-task of Perception-guided Crossmodal Entailment (PCE) which attempts to predict how a test person assesses the coherence of different modalities of a multimodal stimulus (e.g., an image with captions) given the test person’s visual fixation sequence. We collect the first POV learning benchmark data set based on PCE by tracing each test person’s eye movements while this person is judging if the subjectively central visual objects are mentioned in the image’s captions (see Sec. 3).
  2. 2.
    We empirically show how eye tracking sequences can be used for improved individual alignment of POV machine learning models. For this, we explore multiple ways of exploiting the eye tracking data in a variety of machine learning models (see Sec. 4 and 5). To this end, we propose the Perception-Guided Multimodal Transformer (PGMT) which is optimized for PCE (see Sec. 6).
  3. 3.
    Additionally, we assess how multimodal foundation models (i.e., GPT4) perform on POV-learning in zero and few-shot settings (see Sec. 7).
Our findings suggest that augmenting machine learning models with individual perception signals and optimizing for POV-specific objectives is a promising direction of future research. Our PGMT clearly outperforms GPT4 on the PCE task and contributes to the alignment of AI systems on an individual level. We call this new perspective on AI alignment POV Learning.
2 Related work
A variety of research areas are related to the work we present. On a technical level, the field of multi-modal (in our case visual-linguistic) representation learning provides the technical background for our experiments. The novel PCE task that is introduced in this work is related to Visual Entailment tasks and adds a behavioral component to it. Other related work will be covered in the following subsection, where we discuss previous work on eye tracking in psychology and how it relates to POV learning.
2.1 Foundation Models and their Limitations
There are two key techniques on which many recent language models are built on: the attention mechanism [1] and the transformer architecture [24]. Despite their impressive capabilities, there are downsides that come along with models like Gemini [23], GPT4 [15], Meta Llama 3111https://ai.meta.com/blog/meta-llama-3/ and beyond. Firstly, the amount of resources needed to train a modern language model in both time and storage is tremendous. Secondly, the alignment of Large Language Models (LLMs) on an individual level as a POV learning task is an open research question.
2.2 Visual-Linguistic Transformers and Tasks
There exists a variety of recent approaches in Visual-linguistic representation learning. Before the introduction of novel multi modal LLMs like GPT4 [15] and Gemini [23], which are currently the dominant approaches and outperform the traditional models across all benchmarks, they usually combined the BERT [4] architecture with a vision module. Models that fall in this category are VilBert [14], VL-Bert [20] and Visual Bert [11].
Others, like LXMERT [21], Uniter [3], SimvLM [26] and, most recently, OFA [25] use other transformer-based architectures. They all have in common that they are evaluated at least partly on one or more of the benchmark tasks provided by datasets like Visual Genome [10], MS-Coco [12], or Flickr30k [19]. Normally, the tasks are divided into two sub-categories which are briefly described here.
Pre-training tasks:
Masked Language Modeling, Masked Region Modeling, Textual Grounding
Downstream tasks:
Visual Commonsense Reasoning, Visual Question Answering, Natural Language for Visual Reasoning, Region to Phrase Grounding, Visual Entailment
This list is not exhaustive, but contains tasks most commonly referred to in Visual-Linguistic representation learning. The task with the highest similarity to the Perception-guided Crossmodal Entailment task we propose in this work is Visual Entailment [28]. In Visual Entailment, an image text combination functions as a premise/hypothesis pair and the task is to classify whether the hypothesis is entailed in the premise, contradicts the premise, or is neutral with regard to the premise.
2.3 AI Alignment
The goal of AI Alignment is to ensure that AI models behave in a way that overlaps with human values. Due to the ever increasing importance of LLMs and AI models in general, behavior like untruthful responses, hallucinating, or societal biases needs to be mitigated to prevent harmful effects on society [8]. To prevent such harmful responses, originally reinforcement learning from human feedback (RLHF) was introduced in the deployment process of LLMs. This technique relies on human provided feedback, which serves as an alignment signal to prevent LLMs from returning harmful output [16] and has since seen continuous improvements.
However, in this work, we divert from the global population-level view on alignment and instead focus on alignment on the individual level. Human values differ on all levels of granularity in human societies and cultures. The alignment on fine-grained levels down to the individual point-of-view has not been addressed so far. Here, we focus on an alignment signal that reflects individual behavior.
2.4 Machine Learning on Eye Tracking Data
Historically, eye tracking data has been most common as a signal in the psychology of (visual) perception, where it has been used to determine how visual perception works, both physiologically and psychologically [5]. In other lines of psychology related work, eye tracking data is used to to train statistical models for predicting behavior in binary choice settings (cmp. [9] [22] [2]). In contrast to our work, the works described above use directly derived information from the eye tracking sequences (i.e. dwell time, revisits) [18] as input for their statistical models, while we focus on symbolic representations of multi-modal AOIs [^2] In our case, we describe all annotated regions of the stimuli as AOI. These can be both the words in the caption and the visual entities in the image. that we use as input for deep learning models.
In summary, there are important key points in which our work differs from both the line of Visual-Linguistic Transformer research and the line of eye tracking based research in psychology. In contrast to Visual Entailment tasks, we focus on the behavioral aspect of visual perception which we use as individual alignment signal. In Visual Entailment, there is a single population-wide ground truth for a given image/text pair, whereas in the POV learning task, different participants might have opposing evaluations of a given image/text pair, which is why our models have to reflect this aspect in their inferences. On the other side, the models in Psychology do account for individual behavior, but are purely descriptive statistical models (as opposed to our inferential deep learning based approaches), use the data differently (feature engineering on the basis of some properties in the eye tracking sequence, vs. representation learning techniques on our end) and do not account for multi-modality.
Most importantly, no previous work attempts to fuse the two research areas in order to derive individually aligned machine learning models.
3 Perception-guided Crossmodal Entailment
This section describes how the data for both, the perception-guided ML model based on eye tracking data, as well as the content based transformer models were obtained. For both models the identical prediction targets are used, namely the task of Perception-guided Crossmodal Entailment (PCE).
As basis for the eye tracking study, we selected a set of multi-modal documents from the Visual Genome data set [10]. Visual Genome (VG) consists of 108,249 images that are annotated on several layers, from regions to objects, represented in a graph structures. We leverage the underlying scene graph to rank entities depicted in an image by their centrality degree. Based on this ranking, we chose regions that are connected by their contained entities and retrieve the associated caption. We extract up to three regions for each picture which, due to the procedure just described, are connected. The region captions then serve as textual description of the picture. The region descriptions combined with the image constitute the multi-modal document used for further annotation.
3.1 POV Task Specification
To incentivize the test persons we showed the question “Are the central objects in the image mentioned in the caption?” to the study’s participants since it requires crossmodal reasoning and it assesses human perception.
For the eye tracking study 153 stimuli were generated. Errors in the captions and object annotations were corrected. Objects with incorrect VG-annotations were removed. This was documented so that all preprocessing steps can be reconstructed afterwards.
3.2 Eye tracking and Indvidual Assessment Recording
Overall, 109 participants in the age range of 19-61 years participated in the study, which took place in the eye tracking laboratory of the Media Studies department at our institution. The majority of the participants were female (F= 77, 70.64 %; M= 30, 27.52 %), the average age was Mage = 25.4 years. They were recruited via several open calls on university newsletter sites, via online advertisements on social media (Facebook, Instagram), and on printed posters on campus. The experiments were conducted between 5 May and 14 July, 2022. A small participant remuneration was offered (5 € per participant). Participants were mainly students and staff of the University. Participation requirements included basic knowledge of English to ensure they understood the textual stimuli in the experimental task.
A total of 153 text-image-stimuli were selected to be shown to the participants. To ensure that the length of an eye tracking session was not too long to retain the participant’s attention, the participants were divided into three groups: group A (Nparticipants=37; Nstimuli=50), group B (Nparticipants=35; Nstimuli=50) and group C (Nparticipants=37; Nstimuli=53). With exception of the different stimuli, all other experiment settings were the same: a short instruction slide was shown followed by 50-53 blocks consisting of a text-image-stimulus and a following question slide. The blocks of stimulus and question slides were shown in a randomized order. For the study, the experiment was programmed and analyzed in the software iMotions (Version 9.1.5). The total duration of the experimental task was approximately eight minutes (excluding the time for initial calibration).
3.3 Symbolic Fixation Sequence Extraction
From the eye tracking study we obtain the participant-specific fixation sequences for all stimuli. In addition to the information about the participant and stimulus, each fixation contains information about its location (x and y coordinates), its duration, dispersion and the associated annotated region. For instance, Table 1 lists how the sequence in Figure 3 is represented. All fixations are ordered by the time of their occurrence, meaning that they are ordered chronologically.
Figure 3: Fixation Sequence with a negative response.
Note that in this raw data export from iMotions, there are artifacts like in row numbers 2 and 3, where the same fixation is assigned to two different AOIs. In cases like this, we removed the non-specific entry for AOI off, which normally is used to describe a fixation that is on none of the annotated AOIs. The coordinates represent fixation centers calculated by iMotions, which is why they are floating point numbers instead of integers, as one would expect for pixel values.
Table 1: Fixation sequence for Participant EWCX and Stimulus ID 2412873.
Nr. AOI Coordinates Duration
1 vis_wall 865.44, 346.46 100.00
2 txt_wall 709.13, 29.92 158.32
3 off 709.13, 29.92 158.32
4 txt_Rock 658.71, 39.78 191.67
5 txt_wall 760.30, 32.51 258.33
18 vis_ground 976.25, 859.95 166.66
19 vis_zebra 718.85, 823.35 166.66
20 vis_wall 988.11, 431.43 250.00
3.4 Data summary
In summary, the data we obtain consist of the created stimuli (images, annotated AOIs and corresponding captions) and a participant specific fixation sequence for each stimulus. Overall, the acquired data set contains 5,400 unique fixation sequences, where each sequence represents a perception pattern of a participant generated when parsing the stimulus. The fixation sequences contain a total of 148,100 identified fixations, on average 27.42 fixations per stimulus exposure. The distribution of the prediction targets is heavily skewed, with 67.4 % of the samples belonging to the Yes class, 20.1 % to the No class, and 12.5 % to the Unclear class. The data set (including raw fixation data, stimuli, caption texts and annotated regions) and the code for the experiments is available on GitHub. [^3] https://github.com/siwer/Crossmodal-Entailment
4 POV Models and Experimental Design
To test our hypothesis that the evaluation process of stimuli is reflected in the eye tracking data and can act as a feature for improving individual alignment, three different models with focus on different aspects of our data were implemented.
The focus of the LSTM [7] model is on exploring the informational value of the fixation sequence of symbolic components of the multimodal stimulus as alignment signal, without relying on content information about the visual and textual entities in the stimuli.
In contrast to this, the focus of the Transformer model is on content-based information in form of pre-trained feature representations of the stimuli. To enable the Transformer to being able to take into account individual preferences, we also provided a rudimentary participant representation as an alignment signal.
Finally, as a combination of both, we train an Ensemble model that evaluates the predictions of both models jointly.
The task for all models is to predict each participant’s evaluation of the text-image combinations individually. This problem can be formulated as a classification problem with three possible outcomes: 1) yes, the caption mentions the central objects in the image, 2) no, central objects in the image are not mentioned and 3) unclear.
4.1 Technical Setup
The data is divided into training, evaluation and test sets in a random 80/10/10 split, with every split retaining the global class distribution from the data. All the models that are described in the following sections are optimized with AdamW [13] optimizer on a Cross-entropy loss and trained for a maximum of 30 epochs. The final model is chosen as the one with minimum loss on the evaluation split. If not stated otherwise, we use the PyTorch [17] (version 1.11) implementations for all building blocks in our models. We conduct all experiments on a GPU server at our institution.
The hyperparameter ranges that were tested to find the best performing Ensemble and best performing Perception-guided Transformer are stated in table 2. All possible combinations were evaluated. We report accuracy and F1 score as the evaluation metrics to account for the class imbalance in the data. Because of the sparsity of the unclear class, we test all models in two different settings, one with all classes present (3 classes) and one with the unclear class excluded from evaluation.
Table 2: Tested hyperparameters and their ranges.
Parameter Range
Learning Rate 1e-04, 5e-04, 1e-05, 5e-05
FF-Dim 32, 64, 128, 256
Embedding Dim 8, 16, 32
Batch size 16, 64, 128
4.2 Perception-Guided LSTM Model
For this setup, we implement a unidirectional LSTM model that solely receives symbolic representations of the AOIs as a sequence. In our case, we refer to all annotated regions from the eye tracking study as AOIs. If we take the sequence from Table 1 as an example, each row corresponds to one step in the LSTM. In this example, the symbol denoted vis_wall would correspond to the input in the first step t1, txt_wall to t2, and so on until the last element of the sequence vis_wall is reached [^4] Note that we do not add any additional information to the sequence, so that representations are purely symbolic. Before feeding the sequence into the LSTM, we put the one-hot encoded elements through an embedding layer that outputs a dense low-dimensional vector representation of the elements in the AOI sequence. Additionally, we apply the same technique to create embeddings for the participants. As a final step we concatenate each AOI sequence element with the current participant embedding and put it through a feed-forward layer to receive a combined representation, capturing Who is looking at which AOI. This final representation is then used as the input for the LSTM, of which we take the final hidden state and perform the classification by running it through a final feed-forward layer which serves as classification layer. During training, the Embedding layers are also updated, so that the resulting representations are fitted to the task. Overall, there are 837 unique AOI identifiers (i.e. the text representations like ”vis_wall” or ”txt_zebra”) in the dataset. The symbolic AOI representations are not stimulus specific, but rather on a conceptual level, which is why the ”vocabulary” is relatively small. This setting can be seen as using only the individual alignment signal, without the stimulus itself.
4.3 Content-based Transformer Model
The fundamental building block for this baseline is a transformer encoder, which comprises multiple attention heads and layers. Figure 4 shows the baseline model, which can be extended and adapted to different tasks by adding additional layers on top. We use pre-trained BERT [^5] https://huggingface.co/docs/transformers/v4.28.1/en/model_doc/bert [4] embeddings that are obtained from the Huggingface Transformers Package [27] as representations for the linguistic inputs (the captions) and a pre-trained ResNet50666https://pytorch.org/vision/main/models/resnet.html [6] for obtaining image features of the pre-defined regions. The reason for using pre-defined regions instead of applying an object detector as in most state-of-the-art approaches is to ensure a one-to-one mapping between eye tracking data and machine learning input.
Figure 4: Baseline Multimodal Transformer: Textual input is referred to as $w$ and image regions as $r$, respectively. The outputs marked as $cm$ are cross-modal embeddings carry both visual and linguistic features
For the experiments described below, we additionally concatenated a low-dimensional participant representation to the cross-modal classification token. The main reason for this is to function as rudimentary alignment signal. There is also a single feed-forward layer between the ResNet image features and the transformer encoder, which is used for downsizing the image feature dimension to match with the BERT Embedding dimension. For the final decision, the classification token is put through a single feed-forward layer with three output dimensions that correspond to the three possible outcomes. Note that the ResNet and BERT weights are not updated during training. Conceptually, the main advantage of the Transformer architecture is that it can use background knowledge in form of the pre-trained representations of the input modalities to solve the Perception-guided Crossmodal Entailment task.
4.4 Ensemble Model
In addition to the perception-guided approach of the LSTM and the content-based approach of the Transformer model, we introduce a meta-model as a combination of both. In this Ensemble model, we simply combine both of the models described above and use the average of both classifications as a final classification. This has been shown sufficient in our case, where we train the Ensemble parts in a joint fashion and do not require to introduce additional trainable parameters. By training both the LSTM and Transformer model together, we assume that the model learns to differentiate between the cases where one model performs better than the other, thus it decides which of the models returns the better classification in a given case.
5 Experiments I - POV Models based on Fixation Sequences
In a first set of experiments, we investigated how the different kinds of information (content vs. sequence) contribute to solving the PCE task on an individual level. For a fair comparison, all three models are trained using the same hyperparameters that are based on the best-performing Ensemble model. We use 6 attention heads and layers for the transformer, 256 dimensions for the transformer feed-forward network and 32 embedding dimensions (for both the participant and AOI embedding) in the LSTM network. All three models are trained using a batch size of 128 and a learning rate of 0.0001. We report the prediction accuracy and the F1 score (averaged over all classes) on the test split as evaluation metrics for all conducted experiments. Since the rare unclear class is problematic for purely sequential models, we also report Accuracy and F1 scores that excludes samples with this class (’2 Class’ setting). In addition to our models, we report a naive baseline that only predicts the most frequent class. Table 3 shows that the combination of sequential and contextual information in an ensemble leads to the best results.
Table 3: Experimental Results for the PCE task.
Setting Metric LSTM Transformer Ensemble Naive
3 Class Accuracy .7466 .7238 .7728 .6715
F1 .4849 .5240 .5628 .2678
2 Class Accuracy .8494 .8611 .8742 .7639
F1 .7767 .7879 .8112 .4330
These results clearly confirm our hypothesis that the fixation sequence contains information that can be used as alignment signal in a POV learning task. Additionally, they show that the combination of background and sequential information in an Ensemble model leads to the best results.
To test the importance of providing a participant representation to aid the alignment we investigate how much the individual participant embeddings (see Sec. 4.2) of our models contribute to their overall performance. We want to find out whether it is sufficient to use individual behavior (the fixations sequences) for alignment, or if the models are improved by a dedicated representation of an individual. For this experiment, we use the same hyperparameters that were used to achieve the results in table 3. The only difference is that here we removed all individual participant’s representations from the models. We report the same evaluation metrics as before and show the relative change in percent.
Table 4: Ablation Results for the PCE task without Participant Embedding.
Setting Metric LSTM Transformer Ensemble
3 Class Accuracy .6944 .7189 .7271
(-6.9 %) (-0.6 %) (-5.9 %)
F1 .3849 .5146 .4956
(-20.6 %) (-1.8 %) (-11.9 %)
2 Class Accuracy .7899 .8254 .8242
(-7.0 %) (-4.1 %) (-5.7 %)
F1 .6144 .7742 .7429
(-20.9 %) (-1.7 %) (-8.4 %)
It becomes apparent that even our simple participant representation that we incorporate into our models contributes positively to individual alignment.
6 Experiments II - Perception-Guided Multimodal Transformer
We propose a new way to combine content and sequential information by adapting a Transformer model. The novel idea is to directly apply the fixation sequences in form of transition matrices by adding them to the attention weights in the multimodal Transformer model. The intention behind this is to find an alternative way of combining the contextual and the sequential data in one model. We call this a Perception-Guided Multimodal Transformer (PGMT). An advantage of this procedure is that it does not require an increased number of trainable parameters when compared to the vanilla transformer approach. Table 5 shows the transition matrix derived from the first 5 rows of table 1.
Table 5: Transition Matrix (partial) for Participant EWCX and Stimulus ID 2412873.
vis_wall txt_wall off txt_rock
vis_wall 0 1 0 0
txt_wall 0 0 1 0
off 0 0 0 1
txt_rock 0 1 0 0
The attention manipulation process that we chose is straight-forward. We use the src_mask parameter in the PyTorch Transformer Encoder to add the transition matrix to the attention weights. [^7] Note that although the parameter is named src_mask, the documentation states that If a FloatTensor is provided, it will be added to the attention weight The attention weight refers to the term in brackets in equation 1 [24].
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
Due to the sparsity of the matrices, and the final softmax application that happens after adding the matrices to the weights, it is beneficial to scale the weights to ensure that the perception information is not vanishing during calculations. This is enabled by the manual tuning parameter $\lambda$. We found it also beneficial to transpose the matrix along the diagonal and add the result to the original matrix. Equation 2 shows this amplification function. An example based on the matrix in table 5 is shown below.
[…]
When we input the matrix from table 5 into equation 2 and set the weight variable $\lambda$ to 5, we get Equation 3:
[…]
This operation reduces the sparsity of the matrix and in combination with the upscaling lead to the best results. When evaluated on the test data, the PGMT achieves an on-par performance with the Ensemble model in both settings. Table 6 shows that the approach we chose leads to competitive results, with the advantage of needing less trainable parameters.
Table 6: Comparison of the Perception-Guided Multimodal Transformer with the Ensemble model.
Setting Model Accuracy F1
3 Classes PGMT .7761 .5294
Ensemble .7728 .5628
2 Classes PGMT .8715 .8196
Ensemble .8742 .8112
In this setup, we used 128 feed-forward dimensions in the transformer, a batch size of 16 and a learning rate of 0.0001. In this setting, the transition matrices are weighted by a $\lambda$ of 500, which we empirically determined to be most effective in this scenario.
7 Experiments III - POV Learning with GPT4
With GPT4 being the state-of-the-art Mulitmodal Large Language Model we investigate how well it performs on POV tasks like PCE and if it is able to utilize perception information to improve individual alignment through few-shot learning.
In our GPT4 experiments, we investigate i) how well a foundation model like GPT4 is capable of solving our PCE task and ii) if few-shot incontext learning can improve its performance. As before, the task for the model is to predict how the individual participants evaluated the question ’Does the caption mention the central entities in the image?’, given a certain stimulus. As with the experiments before, there are three possible outcomes: Yes, No, Unclear. The only difference is, that we now can prompt the model directly, without optimization on this task. The different setups we investigated vary in the amount of participant-specific information that is provided in the prompt. The data and the training and test splits are the same as before. The different setups are as follows:
Zero-shot:
No subject-specific information is provided. The model is tasked to evaluate the question based only on the stimulus, completely ignoring individual assessments of stimuli.
Fixations:
In addition to the stimulus itself, the subject-specific fixation sequence for the current stimulus is provided in the prompt.
One-shot:
This setup provides an additional training sample as an example of the current individual’s behavior and adds it to the system prompt. Thus, each prompt additionally contains a stimulus, a corresponding fixation sequence and the individual response.
Two-shot:
In this setup, an additional second training sample is provided for each prompt. The second sample was chosen to contain a different individual response than the first sample to provide both, a positive and a negative sample.
As before, we report Accuracy and the F1 score in both the 2 class (Unclear class removed for evaluation) and the 3 class setting as performance metrics for the experiments. The experiment was realised with requests to the OpenAI API.
Table 7: Experimental Results for GPT4 alignment.
Setting Metric Zero-shot Fixations One-shot Two-shot
3 Class Accuracy .6568 .6732 .6584 .6683
F1 .3220 .3265 .3132 .3250
2 Class Accuracy .7472 .7657 .7490 .7630
F1 .5156 .5225 .5018 .5210
The are two noteworthy outcomes regarding our GPT4 research questions: Concerning ii) it appears that GPT4 cannot benefit from few-shot learning for the PCE task. At least not in the way we provided fixation information and demonstrations of individual behavior in the prompt. The stimulus might be too complex, for only 1 or 2 training samples. Still, incontext learning with the fixation sequence only improves the performance. Concerning i) we found that overall GPT4’s performance in PCE is considerably and consistently lower than all approaches presented before. This suggests that GPT4 has not been instruction-tuned on a task similar to PCE and that it struggles with this task in general and with individual alignment in particular. Obviously, further research is needed to assess if post-training helps for POV tasks.
8 Conclusions
In this paper we introduce POV Learning. By demonstrating the potential of taking a subject’s point-of-view and incorporating individual characteristics into machine learning models, we can improve predictive performance on an individual level. We also showed that state-of-the-art foundation models like GPT4 struggle with alignment on an individual level on our newly introduced Perception-guided Crossmodal Entailment (PCE) task.
In three different experiments, we explored different ways of using human perception sequences as alignment signals for POV-learning tasks. To determine the relevance of the information contained in a fixation sequence, we implemented an LSTM model that processes a purely symbolic representation of the AOIs in the order of the fixations. This information alone leads to solid results in the PCE task, with an accuracy value of .7466, compared to .6715 for a naive model and .7238 for a content-only transformer model. Combining the LSTM with a Transformer that processes additional contextual information in form of pre-trained features of the text and image batches of the multimodal stimulus, the resulting Ensemble model achieves a further improvement up to an Accuracy of .7728. From this experiment we conclude that the usage of human perception data leads to improved individual alignment in the PCE task.
Additionally, we propose the Perception-Guided Multimodal Transformer (PGMT), which manipulates the attention weights according to sequence information, which improves the Accuracy to .7761. In comparison to the Ensemble model, this approach has the advantage that the extra parameters of the LSTM model can be avoided while the number of parameters in the Transformer is the same.
With an Accuracy of only .6732, GPT4 is clearly outperformed by the PGMT and struggles with both, the PCE tasks and in-context learning in this setting.
In addition, we conduct an ablation study of the first experiments to determine how much the vector representations of the individual participants contribute to the performance. This ablation shows that even a rudimentary representation improves alignment by a considerable margin. Without the individual representations, the model performance decreases up to 20 percent. This further demonstrates the importance of individual alignment signals.
In summary, we presented a new benchmark data set for the novel Perception-guided Crossmodal Entailment task PCE for POV Learning and proposed the PGMT model which can be used to explore the possibilities of aligning human-like perception with transformer models.
9 Future work
The most obvious direction for future work is towards generalization of the approaches presented in this paper to other tasks and models. In a first direction, a reliable and cheaper (in comparison to conducting an eye tracking study) method for capturing human fixation sequences or other cues for individual alignment has to be developed. Without that, the benefits of perception-guided POV models cannot be tested in settings where multimodal stimuli are observed “in the wild”. More generally, identifying and obtaining valuable alignment signals is a future challenge for further research in the field of individual alignment and POV learning.
Another interesting future study should be concerned with the performance of GPT4. Apparently, the proposed task and few-shot prompts did not result in a convincing performance. Additional approaches to individualize MLLM’s inferences via fine-tuning or persona-prompts need to be investigated.